# Data-Driven Chance-Constrained Multiple-Choice Knapsack Problem: Model, Algorithms, and Applications

### v1

@@ VI Computational Study
Experiments are carefully designed to evaluate the performance of DDALS in solving DDCCMCKP. We also conducted ablation studies, which involves disabling individual components of DDALS, to assess the contribution of the components integrated in DDALS. The source code of algorithms and the benchmark sets, are available at Github  [^1] Github: github.com/eddylxf23/DDCCMCKP. All experiments were run in python on an Intel(R) Xeon(R) CPU E5-2699A v4(2.40GHz,252G RAM) with Ubuntu 16.04 LST.
VI-A Algorithm Settings
We set different parameter combinations of weighted parameter $\lambda$ and iteration times $t$ to evaluate the performance on small-scale and large-scale DDCCMCKP respectively. The performance is evaluated by the cost of final solution when the algorithm reaches the maximum of iteration, and the experiment was repeated for 10 times and take the average value. The value of $\lambda$ is tested from 1 to 10. As show in Fig. 5, there are several points worth noting. In the experiments of LAB-ss2, when $T_{\max}=18$, the current solution of the algorithm remains unchanged. The results show that the lowest cost was obtained when $\lambda=0$. However, when we validate the output solution with real confidence level, we find that the solution was infeasible, which means that the algorithm could not find feasible solutions based on the samples without considering the standard deviation. When $T_{\max}=26$, different lambdas do not affect the algorithm’s convergence to the lowest cost solution. In the experiments of LAB-ls2, when $T_{\max}=25$, the algorithm converges very slowly when we set $\lambda=0$, while other settings can converge quickly. However, when $T_{\max}=32$, different lambda values have little effect on the algorithm. That’s because the constraints are relatively loose, and feasible solutions account for a large proportion of the entire solution space, making the effect of standard deviation less significant. In general, it is necessary to set a non-zero value for $\lambda$, and our algorithm is robust to different $\lambda$, with $t=30$ being appropriate.
@@ VI-A Algorithm Settings
Figure 5: Performance analysis of different $\lambda$. As the algorithm iterates, lower cost can be achieved. The results show that the surrogate weight is necessary for DDALS while a non-zero $\lambda$ is necessary to the algorithm.
VI-B Performance of Evaluation Methods
Since each candidate solution can be seen as an unknown random distribution, concentration inequalities can serve as a reference to evaluate the quality of these candidate solutions. We utilize two kinds of concentration inequality as the lower bounds: Bernstein’s Inequality and $K$-order Hoeffding’s Inequality, which have been widely used in the literature for evaluating the performance of optimization algorithms (Please refer to Appendix 1 for detailed formulas).
TABLE III: Evaluation for LAB-ss2
@@ VI-B Performance of Evaluation Methods
We test exact method on LAB-ss2, and calculate the real confidence level for solution $S_{1},S_{2},S_{3},S_{4},S_{5}$ through the original generator, denoted as ‘Real’. ‘B’s’ and ‘H’s’ represent Bernstein’s and Hoeffding’s inequality, respectively. From Table III, we found a substantial gap between the evaluation outcome and the real confidence level, indicating that the real distribution of solution of DDCCMCKP is far different from Gaussian distribution. Results show that the estimate of the exact method is closest to Real, and the estimation error becomes smaller as the real confidence level increases.
TABLE IV: Evaluation for APP-ls1
@@
Next, we test Monte Carlo evaluation on APP-ls1 in Table IV. ‘MC(10e2)’ represents that the evaluation is carried by 10e2 times of Monte Carlo estimation. The result is the average of 10 times’ results. ‘81.1 $\pm$ 3.78’ represents that the mean of estimated confidence level is 81.1% and the standard deviation is 3.78%. The result shows that the estimation error of $210e$ 2 times has the largest fluctuations and $510e$ 5 tends to be close to the real confidence value, but still not enough to meet the requirement for larger scale of problem and higher confidence level. Considering the trade-off between accuracy and speed, we set $610e$ 6 as the Monte Carlo simulation times.
VI-C Results on Small-Scale and Large-Scale DDCCMCKP
To evaluate the effectiveness of DDALS, we compare DDALS with several baseline methods on various benchmark sets from the real-world application. The baseline methods include Greedy, GA and EDA. Greedy typically works well on problems where the locally optimal choice leads to an overall optimal solution. GA performs well in solving different optimization problems [40]. While EDA achieves great success in both combinatorial and continuous optimization [41].
For fairness, we first run DDALS then record the value of ET. Next, we let the Greedy, GA, and EDA keep running until convergence or the ET of which exceeds that of DDALS, then output the current solution as the result. The population size is set to 10 and the number of dominant individuals is 6 for both GA and EDA. For GA, the crossover probability of parents is set to 0.1 and the probability of variation is $1/m$. The real confidence level is calculated by using the original generators to generate $710e$ 7 data. Note that a solution is regarded as feasible when its real confidence level is larger than $0.985$, which means that we allow for an error margin of $0.5\%$.
@@ VI-C Results on Small-Scale and Large-Scale DDCCMCKP
TABLE V: Performance on benchmarks of LAB. The highest average real confidence level and the lowest cost are highlighted through bolding and being underlined.
Benchmark small $T_{\max}$ large $T_{\max}$
C ET ECL RCL FSR C ET ECL RCL FSR
ss3-Greedy 24.4$\pm$0.0 2.0$\pm$0.0 100$\pm$0 99.9$\pm$0.0 1 21.8$\pm$0.0 3.0$\pm$0.0 100$\pm$0 99.9$\pm$0.0 1
ss3-DDALS(O) 21.4$\pm$0.0 291.9$\pm$6.2 99.9$\pm$0 99.5$\pm$0.0 1 12.6$\pm$0.0 254.4$\pm$4.7 100$\pm$0 99.9$\pm$0 1
ss3-DDALS(V1) 21.4$\pm$0.0 291.9$\pm$6.2 99.9$\pm$0 99.5$\pm$0.0 1 12.6$\pm$0.0 254.4$\pm$4.7 100$\pm$0 99.9$\pm$0 1
ss3-DDALS(V2) 21.4$\pm$0.0 291.9$\pm$6.2 99.9$\pm$0 99.5$\pm$0.0 1 12.6$\pm$0.0 254.4$\pm$4.7 100$\pm$0 99.9$\pm$0 1
@@ C ET ECL RCL FSR C ET ECL RCL FSR
The results on LAB benchmarks are presented in Table V. ‘ss3-Greedy’ in the left of table indicates the result of Greedy tested on benchmark ss3. ‘DDALS(O)’, ‘DDALS(V1)’, ‘DDALS(V2)’ and ‘DDALS(V3)’ represents the original, first, second and third version of DDALS. ‘small $T_{\max}$’ and ‘large $T_{\max}$’ are corresponding to small and large $T_{\max}$ in Table I and Table II. Five indicators are utilized to measure the performance. The cost of solution is denoted as ‘C’. The evaluation times ‘ET’ is utilized to compare the convergence speed of algorithms. ‘ECL’ indicates estimated confidence level, while ‘RCL’ is calculated by real generator. ‘FSR’ represents the ratio of actual feasible solution in the output results of the 10 experiments. All results are based on statistical analysis from 10 repetitions of the experiment, denoted as ‘mean $\pm$ standard deviation’. Due to space limitation, parts of results are included in the main text, and the complete results are available in Appendix 5.

### v2

@@ VI Computational Study
In the experiments, we first analyze the sensitivity of different parameter settings of DDALS. Secondly, we adapt the existing methods, GDSO and DRO, to address our problem and compare with DDALS. Thirdly, we further conduct experiments to compare and analyze various commonly used search methods, such as greedy, estimation of distribution algorithm (EDA), and genetic algorithm (GA), combined with solution evaluation approaches against DDALS. Furthermore, we analyze the impact of the integrated acceleration mechanism, evaluating how it enhances the overall efficiency. Finally, we conduct ablation studies, which involves disabling individual components of DDALS, to assess the contribution of the components integrated in DDALS. The source code, benchmark sets and supplementary materials are available online  [^1] Github: github.com/eddylxf23/CCMCKP. All experiments were executed in python on an Intel(R) Xeon(R) CPU E5-2699A v4(2.40GHz,252G RAM) with Ubuntu 16.04 LST.
VI-A Sensitivity Analysis and Parameter Setting
Various parameter combinations of the weighted parameter $\lambda$ and iteration times $t$ to assess the algorithm’s performance are employed. The performance metric is the cost of the final solution when LSS reaches its maximum iteration times. The experiments are repeated 10 times, and the average value is computed. The parameter $\lambda$ is varied from 1 to 10. The results are shown in Figure 5. In the experiments of LAB-ss2, when $T_{\max}=18$, the current solution remains unchanged. The results show that the lowest cost was obtained when $\lambda=0$. However, it is revealed that the solution is infeasible, indicating that the algorithm could not find feasible solutions without considering the standard deviation. For $T_{\max}=26$, different values of $\lambda$ had no impact on the algorithm’s convergence to the lowest cost solution. In the experiments of LAB-ls2, with $T_{\max}=25$, the algorithm exhibited slow convergence when $\lambda=0$, while other settings allowed for quicker convergence. When $T_{\max}=32$, different $\lambda$ settings had small impact on the performance since the intensity of chance constraints are low, making the effect of standard deviation less significant. In summary, it is necessary to set a non-zero value for $\lambda$, and our algorithm exhibits robustness across different $\lambda$ values. Meanwhile, $t=30$ is an appropriate choice for iteration times.
@@ VI-A Sensitivity Analysis and Parameter Setting
Figure 5: Performance analysis of different $\lambda$. The results show that the surrogate weight with a non-zero $\lambda$ is necessary for DDALS.
TABLE II: Evaluation on LAB-ss2
@@
TABLE III: Evaluation on APP-ls1
@@
In order to assess the performance of the exact evaluation approach, we calculate the real confidence level for specific solutions (denoted as “Real”) of LAB-ss2, namely $S_{1},S_{2},S_{3},S_{4},S_{5}$ through the original generator. Results in Table II show that the estimation error becomes smaller as the real confidence level increases. Additionally, we conduct Monte Carlo evaluation on APP-ls1 as shown in Table III. “MC(10e2)” represents that the simulation times is set 10e2. The experimental results are the averages of 10 trials. “81.1 $\pm$ 3.78” represents that the mean of estimated confidence level is 81.1% and the standard deviation is 3.78%. The results show that the estimation error of $210e$ 2 times has the largest fluctuations and $510e$ 5 tends to be close to the real confidence level, but still not enough to meet the requirement for larger problem scale and higher confidence level. Considering the trade-off between accuracy and speed, we set $610e$ 6 as the Monte Carlo simulation times.
VI-B Comparison of DDALS with GDSO and DRO
We adapt two existing algorithms, named GDSO [19] and DRO [13] respectively, to solve our problem. GDSO assumes that the distributions of random variable follow Gaussian distributions. Thus, we apply a statistical fitting method to sample data to obtain the mean and variance of item weights. DRO does not make any assumption about the distributions of random variables. Instead, DRO constructs ambiguity sets (such as ellipsoid sets) based on sample data and transform the problem into a deterministic mixed integer linear programming which can be solved by corresponding solvers. Details of the adaptation of these methods are available in the online supplementary.
In the first comparison shown in Table IV, we find that DRO can not find feasible solutions in the LAB-ss2 instance. GDSO also fail to find feasible solutions when $T_{max}=14$. In the second comparison in Table V, we increase the sample size to 500. The result reveals that GDSO could solve the problem normally, while DRO again fail to find feasible solutions when $T_{max}$ is less than or equal to 20. In the third comparison in Table VI, we change the distributions of the random variables to be uniform distribution, keeping sample size 500. GDSO starts to fail to find feasible solutions when $T_{max}$ is less than 35, while DRO can not find any feasible solutions. Actually, the difference between the uniform distribution and the Gaussian distribution is greater than the difference between the mixed distribution and the Gaussian distribution in the LAB. Under the same intensity of chance constraints, the performance of GDSO gets worse. Although GDSO and DRO have the advantage of short solution times for small-scale problems, as the problem scale increases, their running times show explosive growth. The time consumed when the problem scale is greater than $10\times 20$ far exceeds the running time of DDALS, as shown in Table VII.
TABLE IV: DDALS vs GDSO vs DRO comparison 1
$T_{\max}$ Brute Force DDALS GDSO DRO
Time(s) Cost Time(s) Cost Time(s) Cost Time(s) Cost
27 3.81 8.07 0.99 8.07 0.03 8.07 - fail
23 3.86 10.05 1.19 10.05 0.06 10.05 - fail
18 3.94 10.36 1.69 10.36 0.04 10.36 - fail
16 3.73 14 1.83 14 0.06 14 - fail
14 3.60 19.41 2.1 19.41 - fail - fail
TABLE V: DDALS vs GDSO vs DRO comparison 2
$T_{\max}$ Brute Force DDALS GDSO DRO
Time(s) Cost Time(s) Cost Time(s) Cost Time(s) Cost
30 3.46 11.55 0.99 11.55 0.13 11.55 36.5 15.4
28 3.44 11.55 1.1 11.55 0.13 11.55 22.4 27.1
20 3.44 15.4 1.7 15.4 0.21 15.4 - fail
15 3.40 18.91 2.2 18.91 0.16 18.91 - fail
10 3.49 26.32 2.6 26.32 0.16 26.32 - fail
TABLE VI: DDALS vs GDSO vs DRO comparison 3
$T_{\max}$ Brute Force DDALS GDSO DRO
Time(s) Cost Time(s) Cost Time(s) Cost Time(s) Cost
37 3.44 4.17 1.0 4.17 0.13 4.17 - fail
36 3.53 4.17 1.0 4.17 0.13 4.17 - fail
35.5 3.49 4.17 1.1 4.17 0.13 4.17 - fail
34.5 3.48 4.35 1.7 4.35 - fail - fail
TABLE VII: DDALS vs GDSO vs DRO comparison 4
Instance DDALS GDSO DRO
Time(s) Cost Time(s) Cost Time(s) Cost
LAB-ss1 0.75 5.91 0.07 5.91 0.35 7.1
LAB-ss2 0.82 8.07 0.07 8.07 0.31 8.07
LAB-ss3 1.10 10.06 0.08 10.06 0.42 10.06
LAB-ss4 1.36 9.22 0.09 9.22 1.43 10.47
LAB-ls1 7.73 26.65 2.13 26.68 $>$3600
LAB-ls2 7.95 17.3 48.37 17.3 $>$3600
LAB-ls3 60.98 55.67 $>$3600 $>$3600
LAB-ls4 194.41 97.03 $>$3600 $>$3600
The comparison results are summarized as follows:
  • GDSO fails to find any feasible solution under high intensity of chance constraint. Its runtime increases rapidly with the growth of problem scale. As the difference between the real distribution and the Gaussian distribution increases, its performance gets worse.
  • DRO fails to find feasible solutions with insufficient sample size ($\leq 500$) and high intensity of chance constraint. Its runtime exponentially increases with problem scale.
  • Our proposed algorithm, DDALS, proves effective across all these scenarios.
VI-C Testing Results on LAB and APP
To evaluate the effectiveness of DDALS, we conduct experiments to compare and analyze various commonly used search methods, named Greedy, GA, EDA, each combined with evaluation approaches. Greedy is particularly effective in scenarios where locally optimal choices contribute to globally optimal solutions. GA demonstrate proficiency across various combinatorial optimization problems [38, 39, 40], and EDA exhibits great success in both combinatorial and continuous optimization [41].
For equitable comparisons, we initiate the execution of DDALS, recording its evaluation times. Subsequently, Greedy, GA, and EDA start execution until convergence or their respective ET surpasses that of DDALS. The population sizes of GA and EDA are both set to 10, with 6 dominant individuals. The crossover probability for parents of GA is set to 0.1 and a variation probability of $1/m$. The real confidence level (RCL) is computed by generating $710e$ 7 data using the real generators. In our experiments, $P_{0}$ is set to 0.99. A solution is considered feasible if its real confidence level exceeds $0.985$, allowing for a 0.5% error margin.
@@ VI-C Testing Results on LAB and APP
TABLE VIII: Performance on the instances of LAB. The highest average real confidence level and the lowest cost are highlighted through bolding and being underlined.
Instance small $T_{\max}$ large $T_{\max}$
C ET ECL RCL FSR Time C ET ECL RCL FSR Time
ss3-Greedy 21.4$\pm$0.0 3.0$\pm$0.0 99.9$\pm$0.0 99.5$\pm$0.0 1 0.2$\pm$0.0 17.4$\pm$0.0 4.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0 1 0.4$\pm$0.0
ss3-GA 22.4$\pm$1.4 306.0$\pm$3.7 99.9$\pm$0.0 99.7$\pm$0.2 1 31.7$\pm$1.9 14.2$\pm$2.4 263.5$\pm$2.9 99.7$\pm$0.4 99.2$\pm$1.3 0.73 37.9$\pm$1.1
ss3-EDA 24.8$\pm$1.9 305.3$\pm$3.0 99.9$\pm$0.1 99.8$\pm$0.5 0.93 44.1$\pm$3.0 13.7$\pm$1.3 264.5$\pm$2.8 99.9$\pm$0.0 99.9$\pm$0.1 1 42.8$\pm$3.4
ss3-DDALS(O) 21.4$\pm$0.0 288.5$\pm$9.3 99.9$\pm$0.0 99.5$\pm$0.0 1 9.6$\pm$0.5 12.6$\pm$0.0 254.4$\pm$4.7 100.0$\pm$0.0 99.9$\pm$0.0 1 14.7$\pm$0.9
ss3-DDALS(V1) 21.4$\pm$0.0 288.5$\pm$9.3 99.9$\pm$0.0 99.5$\pm$0.0 1 9.6$\pm$0.5 12.6$\pm$0.0 254.4$\pm$4.7 100.0$\pm$0.0 99.9$\pm$0.0 1 14.7$\pm$0.9
ss3-DDALS(V2) 21.4$\pm$0.0 288.5$\pm$9.3 99.9$\pm$0.0 99.5$\pm$0.0 1 9.6$\pm$0.5 12.6$\pm$0.0 254.4$\pm$4.7 100.0$\pm$0.0 99.9$\pm$0.0 1 14.7$\pm$0.9
ss3-DDALS(V3) 23.9$\pm$1.7 284 99.8$\pm$0.2 99.1$\pm$0.8 0.75 9.1 15.5$\pm$1.7 264 99.9$\pm$0.3 99.5$\pm$0.9 0.9 14.9
ss4-Greedy 30.1$\pm$0.0 2.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1 0.1$\pm$0.0 14.9$\pm$0.0 5.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1 0.3$\pm$0.0
ss4-GA 23.1$\pm$4.4 773.5$\pm$8.0 99.9$\pm$0.2 99.7$\pm$0.4 1 65.6$\pm$2.5 16.4$\pm$3.0 559.3$\pm$8.8 99.8$\pm$0.3 99.4$\pm$0.6 0.9 60.9$\pm$3.1
ss4-EDA 17.4$\pm$0.0 764.5$\pm$3.0 99.9$\pm$0.0 99.9$\pm$0.0 1 98.8$\pm$1.7 13.4$\pm$1.0 554.3$\pm$2.6 99.8$\pm$0.2 99.1$\pm$0.41 0.93 83.5$\pm$3.3
ss4-DDALS(O) 17.4$\pm$0.0 757.9$\pm$31.9 99.9$\pm$0.0 99.9$\pm$0.0 1 21.4$\pm$1.0 11.2$\pm$0.0 538.8$\pm$31.9 99.2$\pm$0.0 99.0$\pm$0.0 1 23.6$\pm$1.1
ss4-DDALS(V1) 17.4$\pm$0.0 757.9$\pm$31.9 99.9$\pm$0.0 99.9$\pm$0.0 1 21.4$\pm$1.0 12.3$\pm$0.7 538.8$\pm$31.9 99.6$\pm$0.1 98.9$\pm$0.2 1 23.6$\pm$1.1
ss4-DDALS(V2) 17.4$\pm$0.0 757.9$\pm$31.9 99.9$\pm$0.0 99.9$\pm$0.0 1 21.4$\pm$1.0 14.1$\pm$0.9 538.8$\pm$31.9 99.9$\pm$0.2 99.0$\pm$0.1 1 23.6$\pm$1.1
ss4-DDALS(V3) 25.3$\pm$4.1 822 99.8$\pm$0.2 99.6$\pm$0.5 1 22.9 14.3$\pm$1.5 514 99.6$\pm$0.3 99.2$\pm$0.7 0.9 24.1
ls4-Greedy 167.5$\pm$0.0 7.0$\pm$0.0 99.2$\pm$0.0 99.1$\pm$0.0 1 3.5$\pm$0.0 167.6$\pm$0.0 8.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1 3.7$\pm$0.2
ls4-GA 124.1$\pm$8.7 11309.8$\pm$7.4 99.1$\pm$0.2 99.1$\pm$0.3 0.97 12310.4$\pm$282.8 107.5$\pm$6.8 8411.4$\pm$9.8 99.1$\pm$0.1 99.0$\pm$0.3 0.93 8437.8$\pm$157.5
ls4-EDA 123.6$\pm$3.6 11305.1$\pm$3.4 99.3$\pm$0.2 99.4$\pm$0.2 1 9498.8$\pm$243.3 116.6$\pm$3.2 8403.9$\pm$3.1 99.6$\pm$0.2 99.7$\pm$0.2 1 7970.2$\pm$162.7
ls4-DDALS(O) 107.1$\pm$1.0 11360.9$\pm$206.6 99.1$\pm$0.1 99.2$\pm$0.1 1 5537.5$\pm$95.9 92.6$\pm$1.1 8483.8$\pm$335.8 99.1$\pm$0.1 99.0$\pm$0.3 0.93 4113.9$\pm$147.9
ls4-DDALS(V1) 109.5$\pm$0.5 11360.9$\pm$206.6 99.7$\pm$0.1 99.7$\pm$0.1 1 5537.5$\pm$95.9 98.8$\pm$2.1 8483.8$\pm$335.8 99.6$\pm$0.1 99.6$\pm$0.1 1 4113.9$\pm$147.9
ls4-DDALS(V2) 108.9$\pm$1.1 11360.9$\pm$206.6 99.6$\pm$0.2 99.6$\pm$0.2 1 5537.5$\pm$95.9 94.4$\pm$2.0 8483.8$\pm$335.8 99.4$\pm$0.2 99.3$\pm$0.3 1 4113.9$\pm$147.9
ls4-DDALS(V3) 110.8$\pm$3.5 11576 99.3$\pm$0.2 99.4$\pm$0.2 1 5651.3 95.1$\pm$2.5 7964 99.1$\pm$0.1 99.3$\pm$0.1 1 3883.5
@@ C ET ECL RCL FSR Time C ET ECL RCL FSR Time
All presented results are derived from a statistical analysis conducted with 30 repetitions, denoted as “mean $\pm$ standard deviation”. Due to space constraints, selected results are included in the main text, while the complete dataset is available in the online supplementary material. Take Table VIII as an example. “ss3-Greedy” represents the results of the Greedy algorithm applied to the LAB-ss3 instance. Variants of the DDALS algorithm are denoted as “DDALS(O)”, “DDALS(V1)”, “DDALS(V2)”, and “DDALS(V3)” for the original, first, second, and third versions, respectively. Categories such as “small $T_{\max}$” and “large $T_{\max}$” correspond to the values of $T_{\max}$ specified in Table I. For example, in experiments on LAB-ss1, “small $T_{\max}$” is set to 11, and “large $T_{\max}$” is set to 14, corresponding to ${11,14}$ in Table I. Five indicators are utilized to measure the performance: “C”represents the cost of the solution, “ET” measures the evaluation times for comparing convergence speeds, “ECL” signifies the estimated confidence level, “RCL” is real confidence level, and “FSR” denotes the ratio of actual feasible solutions in the results of all experiments. “Time” represents the actual runtime of the algorithm on the server, measured in seconds.